OpenStudy (anikate):

(x^2 y^-3)/3y^2 times (y^2)/x^-4

OpenStudy (anikate):

brb afk gotta eat dinner. just leave the steps and ans ry I will be back.

OpenStudy (anonymous):

\[\frac{x^2y^{-3}}{3y^2} \times \frac{y^2}{x^{-4}} \implies \frac{x^2}{3y^5} \times \frac{y^2x^4}{1}\]I just used one fo the laws of exponents: \[a^{-b} = \frac{1}{a^b}\]Can you finish it up?

OpenStudy (anikate):

ummmm

OpenStudy (anonymous):

Hint: Cross cancel

OpenStudy (anikate):

where did you get b from?

OpenStudy (anikate):

nice hint/answer LOLZ

OpenStudy (anonymous):

lol It's just a random variable...it demonstrates the one of the law of exponents.

OpenStudy (anonymous):

Do you get it???????????????????¿???????????????????¿???????????????????¿???????????????????¿

OpenStudy (anikate):

yah

OpenStudy (anonymous):

Alright. What is the final answer then?

OpenStudy (anikate):

x^6/3y^3

OpenStudy (anonymous):

How in the world did I miss this? You are correct XD